OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = quantity x minus 4 over quantity x squared plus 13 x plus 36

OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = quantity x minus 4 over quantity x squared plus 13 x plus 36

OpenStudy (anonymous):

@radar

OpenStudy (superhelp101):

So since it is x that would mean that the denominator is zero Factoring it would look like this (x +4)(x + 9) = 0 ( u set both sides equal to zero) x + 4 = 0 x + 9 = 0 then you just solve :)

OpenStudy (aum):

\[ \Large \frac{x-4}{x^2+13x+36} = \frac{x-4}{(x+4)(x+9)} \] What values of x will make the denominator zero? Those are the vertical asymptotes.

OpenStudy (superhelp101):

@alexandermarrero17 ?